OpenStudy (anonymous):

5(4X + 2 ) = 70 how do you do it ? Thanks

OpenStudy (anonymous):

http://www.wolframalpha.com/bing/?i=5%284x%2b2%29%3d70 this site will provide answer with explanation

OpenStudy (anonymous):

5(4X + 2 ) = 70 20x+10=70 20x=60 x-3

OpenStudy (anonymous):

thats wrong 60/20=3 30x20=600

OpenStudy (anonymous):

its 3.. thanks :)

OpenStudy (anonymous):

5 (4x 2) = 70, 20X+10=70 , 20X=70-10 , 20X=60 , X=60/20 , X= 30